# Thermal and Quantum Phase Transitions of the $\phi^{4}$ Model

### v3

@@ Abstract
The critical behaviour of the classical (thermal) phase transition (CPT) of the Ising model is in the same universality class of the quantum phase transition (QPT) of the $\phi^{4}$ Quantum Field Theory (QFT), this being a paradigmatic example of the relation between a CPT and its quantum counterpart QPT. It is well known that a classical system in $d$ dimension is – for short-range interactions – in the same universality class of the same model in $d-1$ spatial dimension at $T=0$. Nevertheless, to extend the renormalization group (RG) equations of the very same model from $T=0$ to finite $T$ in the usual way resorting to sums on the Matsubara frequencies may lead to subtelties and even to the disappearance of the critical points as extracted from the RG flow, as we comment more in the text. Our goal in this work is to discuss this issue and map out the QPT-CPT diagram focusing as an example on the $\phi^{4}$ QFT, revisiting its relation to the well-known QPT-CPT phase diagram of the Ising model. To this aim, we propose a modification of the usual finite-temperature RG approach by relating the temperature parameter to the running RG scale, $T\equiv k_{T}=\tau k$ where $k_{T}$ is the running cutoff for thermal, and $k$ is for the quantum fluctuations. Once introduced this dimesionless temperature $\tau$, we investigate the consequences on the thermal RG approach for the $\phi^{4}$ model and construct its QPT-CPT phase diagram. Finally, we formulate requirements for the QPT-CPT phase diagram of the $\phi^{4}$ theory based on known properties of the QPT-CPT phase diagram of the Ising model.
@@ I Introduction
A quantum phase transition (QPT) occurs at the quantum critical point at zero temperature where thermal fluctuations are absent. The QPT is understood as a transition between different quantum phases and driven by quantum fluctuations. Contrary to a classical phase transition (CTP) which is a thermal phase transition, the QPT is accessed at absolute zero temperature by varying a physical parameter such as the magnetic field or the pressure. In this work, we refer to this as the ’quantum’ parameter. At finite but very low temperatures, classical and quantum fluctuations compete with each other. In the quantum critical region near the quantum critical point, quantum fluctuations dominate the system behavior and one can study properties of the QPT sachdev . Thus, a so called QPT-CPT phase diagram can be drawn where the ordered and the disordered phases are separated by each other by a critical line which connects the critical temperature and the quantum critical point. Around the classical phase transition, in the classical critical region, the system is governed by thermal fluctuations and this region becomes narrower with decreasing temperatures and converges towards the quantum critical point monotonically.
The Ising model in $d=2$ or $d=3$ dimensions represents an excellent playground to test and to understand the interplay between quantum and classical phase transitions. The quantum Ising model sachdev ; tranverse_ising_1 ; tranverse_ising_2 can be realised by adding a transverse, external magnetic field to the usual Hamiltonian of the classical Ising model. This is the so called transverse field Ising model where the QPT is realised by varying the magnitude of the trannsverse external magnetic field tranverse_ising_1 ; tranverse_ising_2 . The QPT-CPT phase diagram of the Ising model is well-known and its interpretation in terms of effective field theories also deeply investigated sachdev ; shankar . The classical Ising model with vanishing external magnetic field undergoes a second order classical (thermal) phase transition at a finite critical temperature $T_{c}$. The quantum Ising model at zero temperature has a QPT at a critical magnetic field $h_{c}$.
@@ II Main idea
In general, in order to map out the QPT-CPT diagram of the $\phi^{4}$ field theory one has to perform its thermal RG analysis where both quantum and thermal fluctuations are taken into account. This can be done by e.g., using the finite temperature extension of the zero-temperature FRG method. The standard formulation of the FRG method eea_rg is written in Euclidean spacetime in the framework of the zero-temperature QFT. Its generalisation to finite temperature requires the inclusion of the inverse temperature parameter $\beta=1/k_{B}T$ as the upper bound for the imaginary time integral rg_dimful_T_1 ; rg_dimful_T_2 ; finite_T_V_1 ; thermal_rg_phi4 ; thermal_rg_phi4_pressure ; rg_dimful_T_5 ; rg_dimful_T_6 ; rg_dimful_T_8 ; finite_T_V_3 ; rg_dimful_T_10 ; finite_T_V_5 ; thermal_rg_phi4_volume ; rg_dimless_T_1 ; rg_dimless_T_2 ; rg_dimless_T_3 ; rg_dimless_T_4 ; rg_dimless_T_5 ; rg_dimless_T_6 ; rg_dimless_T_7 ; rg_dimless_T_8 ; rg_dimless_T_9 ; rg_dimless_T_10 ; rg_dimless_T_11 .
@@
However, this choice has a drawback, which we think can be useful to be investigated in the context of a QFT discussion of the problem: The explicit $k$-dependence cannot be removed from the dimensionless RG flow equations. For example, the dimensionless thermal FRG equation (3.9) of thermal_rg_phi4 which is given in the so called local potential approximation, has an explicit $k$-dependence and it blows up in the IR limit, i.e., for $k\to 0$ where the attractive fixed points are always given, so it makes no room for non-trivial fixed points such as the Wilson-Fisher (WF) one. As an example, we mention that in Refs. rancon_1 ; rancon_2 ; rancon_3 the quantum $O(N)$ and the Bose-Hubbard model were studied at finite temperatures and it was shown that in order to properly identify the ”quantum WF fixed point” (i.e. at zero-temperature) and the ”classical WF fixed point” (at finite temperatures, where quantum fluctuations are irrelevant), one need to use different dimensionless coupling constants, see more in the following.
@@
where $k_{T}$ is the running cutoff for thermal, and $k$ is for the quantum fluctuations. This choice can be supported by the following argument. The Wilsonian approach requires the rescaling of the upper bound of the imaginary time integral in every blocking step, so, it has to be linked to the running momentum cutoff $k$. However, in the FRG method one has to take the limit $k\to 0$, therefore, the upper bound cannot be considered as the (inverse) temperature but can be used as a running momentum cutoff $T\equiv k_{T}$. Thus, the temperature is related to the dimensionless quantity $\tau$ which is kept constant over the RG flow. A dimensionless (reduced) temperature has been used in the framework of FRG rg_dimless_T_1 ; rg_dimless_T_2 ; rg_dimless_T_3 ; rg_dimless_T_4 ; rg_dimless_T_5 ; rg_dimless_T_6 ; rg_dimless_T_7 ; rg_dimless_T_8 ; rg_dimless_T_9 ; rg_dimless_T_10 ; rg_dimless_T_11 , however, in these works the typical choice is $\tau\equiv\tau_{k}=T/k$ which means that the the dimensionfull parameter $T$ is kept constant over the RG flow, so, $\tau_{k}$ has a trivial (non-vanishing) dependence on the RG scale $k$. For example, in rg_dimless_T_9 at page 42, in Eq.(132) one finds $\partial_{t}\tau=-\tau$ where $t=\ln(k/\Lambda)$ and in rg_dimless_T_11 at the top of the page 12 one can read $\partial_{t}T/k=-T/k$ which result in the trivial RG scale dependence for $\tau_{k}$. Another example is found in rg_dimless_T_7 at page 4 above equation (21), $\tilde{T}(\Lambda^{\star})=1$ where the intermediate momentum scale $\Lambda^{\star}$ is fixed by the running dimensionless temperature $\tilde{T}$ which is identical to $\tau$. Finally, we refer to rancon_3 at page 7 right below equation (50), where one finds $\tilde{T}_{k}=T/(c_{k}k)$. Therefore, while dimensioneless temperature is introduced as a technical tool in rg_dimless_T_1 ; rg_dimless_T_2 ; rg_dimless_T_3 ; rg_dimless_T_4 ; rg_dimless_T_5 ; rg_dimless_T_6 ; rg_dimless_T_7 ; rg_dimless_T_8 ; rg_dimless_T_9 ; rg_dimless_T_10 ; rg_dimless_T_11 , the study of consequences and the determination of physical properties from the constant dimensionless temperature is lacking, to the best of our knowledge. This makes difficult to understand and to identify thermal and quantum phase transitions and their crossovers.
@@
Thermal field theory is a very well-developed framework for finite temperature applications kapusta and it is employed in many-body physics negele ; kapusta which relates a quantum system in $d-1$ (spatial) dimensions at $T=0$ and the corresponding classic system in $d$ dimensions at finite temperature sachdev . It generates the discretisation of the imaginary time-integral in momentum space, i.e., summation on Matsubara frequencies negele ; kapusta . At nonzero temperatures, classical fluctuations with an energy scale of $k_{B}T$ compete with the quantum fluctuations of energy scale $\hbar\omega$ where $\omega$ is the frequency of quantum oscillations. By using natural units, one finds that $\omega$ must be compared to $T$ and in the RG method one has to relate the frequency $\omega$ to the running cutoff $k$ which is used to integrate quantum fluctuations systematically, so one finds $\omega=k$. Thus, one can introduce a dimensionless quantity, $\tau=T/k=T/\omega$ which is used to compare the strength of quantum and thermal fluctuations. If $\tau\lesssim 1$ then quantum fluctuations dominate, if $\tau\gtrsim 1$ then thermal fluctuations dominate. Notice that in the Wilsonian approach fluctuations are taken into account by the successive elimination (integration) of degrees of freedom above these running cutoffs which are chosen to be different in order to make difference between thermal and quantum fluctuations. In other words, for $k_{T}\sim k$ (i.e., for $\tau\sim 1$) one cannot distinguish between thermal and quantum fluctuations.
@@ III Modified thermal RG equation
$k\partial_{k}{V_{k}(\phi)}=\frac{1}{2}\int_{-\infty}^{\infty}\frac{d^{d}p}{(2\pi)^{d}}\,\frac{k\partial_{k}R_{k}}{R_{k}+V^{\prime\prime}_{k}(\phi)}\,,$ (2)
@@
[…]
@@ III.1 Thermal RG with the Litim regulator
In order to proceed further one has to assume that the regulator function is independent of the Matsubara frequency, but othevise the same as for the zero-temperature case. This is the case of the ”frequency independent regulator”. In addition, it is also important to note that once approximations are used, the concrete form of the FRG equation start to depend on the particular choice of the regulator function. There are two types of regulator functions where the momentum integrals in (2) and its finite-temperature counterpart can be performed analytically in LPA. The first one is the Litim regulator Litim2000 , which has the following form at zero-temperature,
@@
The schematic thermal RG flow with $T=\tau\Lambda$ was shown in Fig. 6 of rg_dimful_T_5 and Fig. 4 of rg_dimful_T_6 (which are identical figures) and given here on Fig. 1.
@@
In this work we use the identification $T\equiv k_{T}=\tau k$ where ”$\tau$” plays the role of the (dimensionless) temperature which distinguishes between thermal quantum fluctuations. We suggest a modified thermal RG equation,
@@ IV The zero-temperature limit
${\rm T\to 0:}\hskip 14.22636ptk\partial_{k}V_{k}(\phi)=\frac{2\alpha_{d-1}}{d-1}\frac{k^{d+1}}{2\sqrt{k^{2}+\partial^{2}_{\phi}V_{k}(\phi)}}\,,$ (11)
@@
${\rm T=0:}\hskip 14.22636ptk\partial_{k}V_{k}(\phi)=\frac{2\alpha_{d}}{d}\,\frac{k^{d+2}}{k^{2}+\partial^{2}_{\phi}V_{k}(\phi)}\,,$ (12)
@@ V Thermal RG study of the $\phi^{4}$ model in lower dimensions
$V_{k}(\phi)=\sum_{n=1}^{\rm NCUT}\frac{g_{2n,k}}{(2n)!}\phi^{2n}\hskip 0.0pt\to\hskip 0.0pt\tilde{V}_{k}(\tilde{\phi})=\sum_{n=1}^{\rm NCUT}\frac{\tilde{g}_{2n,k}}{(2n)!}{\tilde{\phi}}^{2n}$ (13)
@@ VI QPT-CPT diagram of the $\phi^{4}$ model
It is known that the classical Ising model is equivalent to the zero-temperature $\phi^{4}$ QFT. Thus, the vertical axis of the he QPT-CPT diagram and the critical temperature of the Ising model can be mapped onto the horizontal axis of the he QPT-CPT diagram and the critical ’quantum’ parameter of the (zero-temperature) $\phi^{4}$ QFT, see the blue lines of Fig. 4. In addition, the thermal critical behavior of the classical Ising model must be identical to the quantum critical behavior of the $\phi^{4}$ QFT.
@@
For vanishing quantum parameter, i.e., $h/J=0$, the Ising model becomes the well-known classical one, which is non-trivial, it has a classical (thermal) phase transition with a finite transition temperature $T_{c}<\infty$. On contrary, for a vanishing ’quantum’ parameter, i.e., for $(\tilde{g}_{4}/|\tilde{g}_{2}|)_{\rm{WF}}=0$, the $\phi^{4}$ model becomes a trivial massive free field theory (with thermal fluctuations), which has no phase transition, so there is no room for any finite transition temperature. On the one hand, the critical temperature of the $\phi^{4}$ model must be zero if $(\tilde{g}_{4}/|\tilde{g}_{2}|)_{\rm{WF}}=0$. On the other hand, the critical line must be a strictly monotonic function, so one expects $T_{c}\to\infty$ for the $\phi^{4}$ in the limit $(\tilde{g}_{4}/|\tilde{g}_{2}|)_{\rm{WF}}\to 0$. Therefore, one expects a discontinuity in the classical limit of the $\phi^{4}$ theory; this is why we use an empty circle at $\tau_{c}=\infty$ on the lower panel of Fig. 4 and in its inset. In addition, one can conclude that the Ising model has a thermal (classical) and a quantum critical behavior, too but the $\phi^{4}$ model has a quantum critical behavior, only.
@@
However, it is well-known that the critical behavior of the Ising model has a discontinuity in the limit $T\to 0$, see the empty circle at $T=0$ (i.e., at $h_{c}$) on the upper panel of Fig. 4. So, critical lines (green lines on the upper and lower panels of Fig. 4) have a discontinuity at their endpoints: for the Ising it happens at $h/J=(h/J)_{c}$ and for the $\phi^{4}$ theory it situates at $\tau=\tau_{c}=\infty$.
@@
Finally let us discuss an interesting property of the infinity temperature limit. We showed, that in the limit $\tau\to\infty$ the WF fixed point tends to the horisontal axis. In other words, the quartic coupling $\tilde{g}_{4}^{\rm WF}$ of the WF fixed point tends to zero for $\tau\to\infty$. However, one can show that their product, $\tau\tilde{g}_{4}^{\rm WF}$ tends to a constant value. Thus, the QPT-CPT diagram where the vertical axis is chosen to be $\tau\tilde{g}_{4}^{\rm WF}$ becomes almost identical to the well-known QPT-CPT diagram of the Ising model, see Fig. 6.

### v4

@@ Abstract
In this paper we discuss and revisit the finite temperature extension of the renormalization group (RG) treatment of $T=0$ field theories, focusing as a case study on the $\phi^{4}$ model. We first discuss the extension of RG equations of the very same model from $T=0$ to finite $T$ in the usual way by resorting to sums on the Matsubara frequencies and fixing the physical temperature parameter $T$. We show that this approach, although useful for a variety of applications, may lead to the disappearance of the critical points as extracted from the RG flow. Since the identification of fixed points is key in the study of classical and quantum phase transitions, we propose a modification of the usual finite-temperature RG approach by relating the temperature parameter to the running RG scale, $T\equiv k_{T}=\tau k$ where $k_{T}$ is the running cutoff for thermal, and $k$ is for the quantum fluctuations. Once introduced this dimensionless temperature $\tau$, we investigate the consequences on the thermal RG approach for the $\phi^{4}$ model and construct its phase diagram. Finally, we formulate requirements for the phase diagram of the $\phi^{4}$ theory based on known properties of the quantum and classical phase diagrams of the Ising model.
@@ I Introduction
A quantum phase transition (QPT) occurs at the quantum critical point at zero temperature where thermal fluctuations are absent. The QPT is understood as a transition between different quantum phases and driven by quantum fluctuations. Contrary to a classical phase transition (CPT) which is a thermal phase transition, the QPT is accessed at absolute zero temperature by varying a physical parameter such as the magnetic field or the pressure. In this work, we refer to this as the ’quantum’ parameter. As a major example, the critical behaviour of the CPT of the Ising model is in the same universality class of the QPT of the $\phi^{4}$ Quantum Field Theory (QFT), this providing a paradigmatic instance of the relation between a CPT and its quantum counterpart QPT. It is well known that a classical system in $d$ dimension is – for short-range interactions – in the same universality class of the same model in d1d-1 spatial dimension at $T=0$.
At a finite but suitably low temperature, classical and quantum fluctuations compete with each other. In the quantum critical region near the quantum critical point, quantum fluctuations dominate the system behaviour and one can study properties of the QPT sachdev . Thus, a so called QPT-CPT phase diagram can be drawn where the ordered and the disordered phases are separated by each other by a critical line which connects the critical temperature and the quantum critical point. Around the classical phase transition, in the classical critical region, the system is governed by thermal fluctuations and this region becomes narrower with decreasing temperatures and converges towards the quantum critical point monotonically.
The Ising model in $d=2$ or $d=3$ dimensions represents an excellent playground to test and to understand the interplay between quantum and classical phase transitions. The quantum Ising model sachdev ; tranverse_ising_1 ; tranverse_ising_2 can be realized by adding a transverse, external magnetic field to the usual Hamiltonian of the classical Ising model. This is the so called transverse field Ising model where the QPT is realized by varying the magnitude of the transverse external magnetic field tranverse_ising_1 ; tranverse_ising_2 . The QPT-CPT phase diagram of the Ising model is well-known and its interpretation in terms of effective field theories also deeply investigated sachdev ; shankar . The classical Ising model with vanishing external magnetic field undergoes a second order classical (thermal) phase transition at a finite critical temperature $T_{c}$. The quantum Ising model at zero temperature has a QPT at a critical magnetic field $h_{c}$.
@@ II Main idea
In general, in order to map out the QPT-CPT diagram of the $\phi^{4}$ field theory one has to perform its thermal RG analysis where both quantum and thermal fluctuations are taken into account. This can be done by e.g., using the finite temperature extension of the zero-temperature FRG method. The standard formulation of the FRG method eea_rg is written in Euclidean spacetime in the framework of the zero-temperature QFT. Its generalization to finite temperature requires the inclusion of the inverse temperature parameter $\beta=1/k_{B}T$ as the upper bound for the imaginary time integral rg_dimful_T_1 ; rg_dimful_T_2 ; finite_T_V_1 ; thermal_rg_phi4 ; thermal_rg_phi4_pressure ; rg_dimful_T_5 ; rg_dimful_T_6 ; rg_dimful_T_8 ; finite_T_V_3 ; rg_dimful_T_10 ; finite_T_V_5 ; thermal_rg_phi4_volume ; rg_dimless_T_1 ; rg_dimless_T_2 ; rg_dimless_T_3 ; rg_dimless_T_4 ; rg_dimless_T_5 ; rg_dimless_T_6 ; rg_dimless_T_7 ; rg_dimless_T_8 ; rg_dimless_T_9 ; rg_dimless_T_10 ; rg_dimless_T_11 .
@@
However, every approach has advantages and disadvantages. In the fixed T approach the dimensionful temperature is well-defined, the IR limit $k\to 0$ can be taken safely but the RG flow equations have no fixed point solutions, since the explicit $k$-dependence cannot be removed from the dimensionless RG flow equations. For example, the dimensionless thermal FRG equation (3.9) of thermal_rg_phi4 which is given in the so called local potential approximation, has an explicit $k$-dependence and it blows up in the IR limit, i.e., for $k\to 0$ where the attractive fixed points are always given, so it makes no room for non-trivial fixed points such as the Wilson-Fisher (WF) one. As an example, we mention that in Refs. rancon_1 ; rancon_2 ; rancon_3 the quantum $O(N)$ and the Bose-Hubbard model were studied at finite temperatures and it was shown that in order to properly identify the ”quantum WF fixed point” (i.e. at zero-temperature) and the ”classical WF fixed point” (at finite temperatures, where quantum fluctuations are irrelevant), one need to use different dimensionless coupling constants, see more in the following.
@@
where $k_{T}$ is the running cutoff for thermal, and $k$ is for the quantum fluctuations. This choice can be supported by the following argument. The Wilsonian approach requires the rescaling of the upper bound of the imaginary time integral in every blocking step, so, it has to be linked to the running momentum cutoff $k$. However, in the FRG method one has to take the limit $k\to 0$, therefore, the upper bound cannot be considered as the (inverse) temperature but can be used as a running momentum cutoff $T\equiv k_{T}$. Thus, the temperature is related to the dimensionless quantity $\tau$ which is kept constant over the RG flow. A dimensionless (reduced) temperature has been used in the framework of FRG rg_dimless_T_1 ; rg_dimless_T_2 ; rg_dimless_T_3 ; rg_dimless_T_4 ; rg_dimless_T_5 ; rg_dimless_T_6 ; rg_dimless_T_7 ; rg_dimless_T_8 ; rg_dimless_T_9 ; rg_dimless_T_10 ; rg_dimless_T_11 , however, in these works the typical choice is $\tau\equiv\tau_{k}=T/k$ which means that the the dimensionful parameter $T$ is kept constant over the RG flow, so, $\tau_{k}$ has a trivial (non-vanishing) dependence on the RG scale $k$. For example, in rg_dimless_T_9 at page 42, in Eq.(132) one finds $\partial_{t}\tau=-\tau$ where $t=\ln(k/\Lambda)$ and in rg_dimless_T_11 at the top of the page 12 one can read $\partial_{t}T/k=-T/k$ which result in the trivial RG scale dependence for $\tau_{k}$. Another example is found in rg_dimless_T_7 at page 4 above equation (21), $\tilde{T}(\Lambda^{\star})=1$ where the intermediate momentum scale $\Lambda^{\star}$ is fixed by the running dimensionless temperature $\tilde{T}$ which is identical to $\tau$. Finally, we refer to rancon_3 at page 7 right below equation (50), where one finds $\tilde{T}_{k}=T/(c_{k}k)$. Therefore, while dimensionless temperature is introduced as a technical tool in rg_dimless_T_1 ; rg_dimless_T_2 ; rg_dimless_T_3 ; rg_dimless_T_4 ; rg_dimless_T_5 ; rg_dimless_T_6 ; rg_dimless_T_7 ; rg_dimless_T_8 ; rg_dimless_T_9 ; rg_dimless_T_10 ; rg_dimless_T_11 , the study of consequences and the determination of physical properties from the constant dimensionless temperature is lacking, to the best of our knowledge. This makes difficult to understand and to identify thermal and quantum phase transitions and their crossovers.
It is important to note that our proposal for fixing $\tau$ = T/k and taking the simultaneous $T,k\to 0$ limit addresses a complementary regime compared to the widely used fixed $T$ approach. In a certain level of approximation the two methods can give qualitatively same results, so one cannot come to the conclusion that previous studies were incorrect. However, it is also important to note that in our fixed $\tau$ approach the RG flow equations have real (and not pseudo) fixed points which have fundamental importance since the determination of critical behaviour and various phases are strongly related to fixed point solutions of RG flow equations.
@@
Thermal field theory is a very well-developed framework for finite temperature applications kapusta and it is employed in many-body physics negele ; kapusta which relates a quantum system in d1d-1 (spatial) dimensions at $T=0$ and the corresponding classic system in $d$ dimensions at finite temperature sachdev . It generates the discretisation of the imaginary time-integral in momentum space, i.e., summation on Matsubara frequencies negele ; kapusta . At non-zero temperatures, classical fluctuations with an energy scale of $k_{B}T$ compete with the quantum fluctuations of energy scale $\hbar\omega$ where $\omega$ is the frequency of quantum oscillations. By using natural units, one finds that $\omega$ must be compared to $T$ and in the RG method one has to relate the frequency $\omega$ to the running cutoff $k$ which is used to integrate quantum fluctuations systematically, so one finds $\omega=k$. Thus, one can introduce a dimensionless quantity, $\tau=T/k=T/\omega$ which is used to compare the strength of quantum and thermal fluctuations. If $\tau\lesssim 1$ then quantum fluctuations dominate, if $\tau\gtrsim 1$ then thermal fluctuations dominate. Notice that in the Wilsonian approach fluctuations are taken into account by the successive elimination (integration) of degrees of freedom above these running cutoffs which are chosen to be different in order to make difference between thermal and quantum fluctuations. In other words, for $k_{T}\sim k$ (i.e., for $\tau\sim 1$) one cannot distinguish between thermal and quantum fluctuations.
@@ III Modified thermal RG equation
$k\partial_{k}{V_{k}(\phi)}=\frac{1}{2}\int_{-\infty}^{\infty}\frac{d^{d}p}{(2\pi)^{d}}\,\frac{k\partial_{k}R_{k}}{p^{2}+R_{k}+V^{\prime\prime}_{k}(\phi)}\,$, (2)
@@
The extension of the zero-temperature FRG method for finite temperatures requires the modification of the imaginary time $\tilde{t}$ integral $\int d^{d}x\to\int_{0}^{\beta}d\tilde{t}\int d^{d-1}x$ where $\beta=1/T$ and $T$ is the temperature parameter. As a next step, the momentum integral with respect to the imaginary time is modified $\int d^{d}p\to T\sum_{\omega_{n}}\int d^{d-1}p$ by using Matsubara frequencies where the summation is performed over these discrete frequencies; for bosonic degrees of freedom $\omega_{n}=2\pi nT$.
@@ III.1 Thermal RG with the Litim regulator
In order to proceed further one has to assume that the regulator function is independent of the Matsubara frequency, but otherwise the same as for the zero-temperature case. This is the case of the ”frequency independent regulator”. In addition, it is also important to note that once approximations are used, the concrete form of the FRG equation start to depend on the particular choice of the regulator function. There are two types of regulator functions where the momentum integrals in (2) and its finite-temperature counterpart can be performed analytically in LPA. The first one is the Litim regulator Litim2000 , which has the following form at zero-temperature,
@@
The schematic thermal RG flow with $T=\tau\Lambda$ was shown in Fig. 6 of rg_dimful_T_5 and Fig. 4 of rg_dimful_T_6 and given here on Fig. 1.
@@
In this work we use the identification $T\equiv k_{T}=\tau k$ where $\tau$” plays the role of the (dimensionless) temperature which distinguishes between thermal quantum fluctuations. We suggest a modified thermal RG equation,
@@ IV The zero-temperature limit
${\rm T\to 0:}\qquad k\partial_{k}V_{k}(\phi)=\frac{2\alpha_{d-1}}{d-1}\frac{k^{d+1}}{2\sqrt{k^{2}+\partial^{2}_{\phi}V_{k}(\phi)}}\,$, (11)
@@
${\rm T=0:}\qquad k\partial_{k}V_{k}(\phi)=\frac{2\alpha_{d}}{d}\,\frac{k^{d+2}}{k^{2}+\partial^{2}_{\phi}V_{k}(\phi)}\,$, (12)
@@ V Thermal RG study of the $\phi^{4}$ model in lower dimensions
$V_{k}(\phi)=\sum_{n=1}^{\rm NCUT}\frac{g_{2n,k}}{(2n)!}\phi^{2n}\thinspace\to\thinspace\tilde{V}_{k}(\tilde{\phi})=\sum_{n=1}^{\rm NCUT}\frac{\tilde{g}_{2n,k}}{(2n)!}{\tilde{\phi}}^{2n}$ (13)
@@ VI QPT-CPT diagram of the $\phi^{4}$ model
It is known that the classical Ising model is equivalent to the zero-temperature $\phi^{4}$ QFT. Thus, the vertical axis of the he QPT-CPT diagram and the critical temperature of the Ising model can be mapped onto the horizontal axis of the he QPT-CPT diagram and the critical ’quantum’ parameter of the (zero-temperature) $\phi^{4}$ QFT, see the blue lines of Fig. 4. In addition, the thermal critical behaviour of the classical Ising model must be identical to the quantum critical behaviour of the $\phi^{4}$ QFT.
@@
For vanishing quantum parameter, i.e., $h/J=0$, the Ising model becomes the well-known classical one, which is non-trivial, it has a classical (thermal) phase transition with a finite transition temperature $T_{c}<\infty$. On contrary, for a vanishing ’quantum’ parameter, i.e., for $(\tilde{g}_{4}/|\tilde{g}_{2}|)_{\rm{WF}}=0$, the $\phi^{4}$ model becomes a trivial massive free field theory (with thermal fluctuations), which has no phase transition, so there is no room for any finite transition temperature. On the one hand, the critical temperature of the $\phi^{4}$ model must be zero if $(\tilde{g}_{4}/|\tilde{g}_{2}|)_{\rm{WF}}=0$. On the other hand, the critical line must be a strictly monotonic function, so one expects $T_{c}\to\infty$ for the $\phi^{4}$ in the limit $(\tilde{g}_{4}/|\tilde{g}_{2}|)_{\rm{WF}}\to 0$. Therefore, one expects a discontinuity in the classical limit of the $\phi^{4}$ theory; this is why we use an empty circle at $\tau_{c}=\infty$ on the lower panel of Fig. 4 and in its inset. In addition, one can conclude that the Ising model has a thermal (classical) and a quantum critical behaviour, too but the $\phi^{4}$ model has a quantum critical behaviour, only.
@@
However, it is well-known that the critical behaviour of the Ising model has a discontinuity in the limit $T\to 0$, see the empty circle at $T=0$ (i.e., at $h_{c}$) on the upper panel of Fig. 4. So, critical lines (green lines on the upper and lower panels of Fig. 4) have a discontinuity at their endpoints: for the Ising it happens at $h/J=(h/J)_{c}$ and for the $\phi^{4}$ theory it situates at $\tau=\tau_{c}=\infty$.
@@
Finally let us discuss an interesting property of the infinity temperature limit. We showed, that in the limit $\tau\to\infty$ the WF fixed point tends to the horizontal axis. In other words, the quartic coupling $\tilde{g}_{4}^{\rm WF}$ of the WF fixed point tends to zero for $\tau\to\infty$. However, one can show that their product, $\tau\tilde{g}_{4}^{\rm WF}$ tends to a constant value. Thus, the QPT-CPT diagram where the vertical axis is chosen to be $\tau\tilde{g}_{4}^{\rm WF}$, see Fig. 6, becomes identical to the well-known QPT-CPT diagram of the Ising model obtained by Monte Carlo simulation, see Fig. 7.
@@
It is important to note that real (and not pseudo) fixed points can only be determined by keeping $\tau=T/k$ fixed over the RG flow in the simultaneous $T,k\to 0$ limit. This provides us the tool to draw the QPT-CPT diagram of a QFT, such as the $\phi^{4}$ model, see Fig. 6. The critical line which separates the broken and the symmetric phases is determined by these (real) fixed points which can be directly compared to simulation results and measurements performed on the corresponding spin model, in our case on the transverse field Ising model, see Fig. 7. If one keeps T constant over the RG flow no such comparison can be done since in this case the thermal FRG has pseudo (and not real) fixed points. So, the direct comparison between the line of real fixed points of quantum-thermal spin models and their thermal QFT counterparts can only be done in the modified thermal FRG method proposed here.
Figure 7: QPT-CPT diagram of the transverse field Ising model obtained by Monte Carlo simulation and taken from qpt_cpt_transverse_field_ising . This can be directly compared to Fig. 6.